Find an equation of the line satisfying the given conditions.
Throung (0,5); m=-(2/3)

Answer by nerdybill(7384)

Find an equation of the line satisfying the given conditions.
Through (0,5); m=-(2/3)
.
Plug into the "point-slope" form:
y-y1 = m(x-x1)
.
y-5 = -(2/3)(x-0)
y-5 = -(2/3)x
y = -(2/3)x + 5 (this is what they're looking for)
The above is the "slope-intercept" form.
Where
slope (m) is -(2/3)
and
y-intercept is at (0,5)